In triangle ABC,the measure of angle A is twice the measure of angle B. The measure of angle C is 30 degrees more than angle A.How many degrees are therein each angle?
Answer by flame8855(424)

A= 2B
C=A+30
A+B+C= 180
2B+B+A+30=180
3B+2B+30=180
5B=150
B=30
A=60
C=90